# FW-Merging: Scaling Model Merging with Frank-Wolfe Optimization

### v2

@@ Abstract
Model merging has emerged as a promising approach for multi-task learning (MTL), offering a data-efficient alternative to conventional fine-tuning. However, with the rapid development of the open-source AI ecosystem and the increasing availability of fine-tuned foundation models, existing model merging methods face two key limitations: (i) They are primarily designed for in-house fine-tuned models, making them less adaptable to diverse model sources with partially unknown model and task information, (ii) They struggle to scale effectively when merging numerous model checkpoints. To address these challenges, we formulate model merging as a constrained optimization problem and introduce a novel approach: Frank-Wolfe Merging (FW-Merging). Inspired by Frank-Wolfe optimization, our approach iteratively selects the most relevant model in the pool to minimize a linear approximation of the objective function and then executes a local merging similar to the Frank-Wolfe update. The objective function is designed to capture the desired behavior of the target-merged model, while the fine-tuned candidate models define the constraint set. More importantly, FW-Merging serves as an orthogonal technique for existing merging methods, seamlessly integrating with them to further enhance accuracy performance. Our experiments show that FW-Merging scales across diverse model sources, remaining stable with 16 irrelevant models and improving by 15.3% with 16 relevant models on 20 CV tasks, while maintaining constant memory overhead—unlike the linear overhead of data-informed merging methods. Compared with the state-of-the-art approaches, FW-Merging surpasses the data-free merging method by 32.8% and outperforms the data-informed Adamerging by 8.39% when merging 20 ViT models. Our code is open-sourced at here.
@@ 1 Introduction
Figure 1: Performance scaling of FW-Merging across CV tasks. (a) demonstrate robustness to irrelevant models, while (b) show improved performance with relevant models. (c) analyzes performance degradation when incorporating a noisy model initialized from a different pre-trained checkpoint. Detailed results and experimental setup are discussed in Section 4.4 and Appendix B.4.
@@ 3.2 Frank-Wolfe Model Merging
We consider the problem of fine-tuning a pre-trained foundation model on new tasks. Given a pre-trained model $\theta_{0}$ and previously fine-tuned models $\{\theta_{1}^{*},\theta_{2}^{*},\cdots,\theta_{n}^{*}\}$, we aim to fine-tune the $(n+1)$-th model $\theta_{n+1}$ on new tasks as a convex combination of the previous models with a) limited high-quality data and b) optimal merging coefficients.
@@
$\min_{\lambda}\ell\Big{(}\sum_{i=1}^{n}\lambda_{i}\theta_{i}^{*},D\Big{)}\quad\text{s.t.}\quad\sum_{i=1}^{n}\lambda_{i}=1,\quad\lambda_{i}\geq 0,$ (5)
where $\lambda=\{\lambda_{1},\ldots,\lambda_{n}\}$ are the merging coefficients, and $\ell$ is the loss function with a multi-task objective on a training dataset $D$ for new tasks.
@@
$\min_{\theta\in\mathcal{M}}\ell(\theta,D),$ (6)
@@
$\min_{\theta\in\mathcal{M}}\ell(\theta,D)$ $=\min_{\lambda\in\Delta^{n}}\ell\Big{(}\sum_{i=1}^{n}\lambda_{i}\theta_{i}^{*},D\Big{)}.$
@@
${\rm LMO}\!\left(\{\theta_{i}^{*}\}_{i=1}^{n},D,\theta_{t}\right)=\operatorname*{arg\,min}_{s\in\rm\{\theta_{1}^{*},\ldots,\theta_{n}^{*}\}}\langle\nabla\ell(\theta_{t},D),s\rangle$ (7)
@@
Input : Initial solution $\theta_{0}$; Fine-tuned checkpoints $\{\theta_{i}^{*}\}^{n}_{i=1}$; Train-set $D$; FW budget $T$.
@@
1:  if $\theta_{0}\notin\mathcal{M}$ then :=$\mathcal{M}:=\mathcal{M}\cup\{\theta_{0}\}$
@@
4:     if $g_{t}:=\langle-\nabla\ell(\theta_{t},D),d_{t}\rangle\leq\epsilon$ then
@@ 3.3 Design choices of the algorithm
In the case of deep learning, the optimization problem is non-convex, additional efforts are needed to better characterize the loss landscape and prevent the LMO from being dominated by one or a few fine-tuned models, which occurs because the linear approximation of $\ell(\theta,D)$ is an inaccurate sketch of the original objective function. Instead of relying on the argmin of linear subproblem, we fetch the top-$k$ vertices of LMO, $\{\tilde{s}_{j}\}_{j=1}^{k}$. A more subtle top-$k$ operation can be performed in a task-wise fashion if the original objective function involves multi-tasks.
@@
The merging function defined in equation (9) maintains feasibility, i.e., the merged model stays in the convex hull $\mathcal{M}$.
@@
Consider $\ell(\theta,D)$ be $L$-smooth over $\mathcal{M}$, which has two constants: ${\rm diam}:=\max_{\theta_{1},\theta_{2}\in\mathcal{M}}\|\theta_{1}-\theta_{2}\|$ be the diameter of $\mathcal{M}$, and ${\rm subopt}:=\ell(\theta_{0},D)-\min_{\theta\in\mathcal{M}}\ell(\theta,D)$ be the global suboptimality. Consider the soft FW algorithm which introduces the following changes to Algorithm 1:
@@
    $\{\lambda_{j}^{*}\}_{j=1}^{k}=\arg\min_{\lambda\in\Delta^{k}}\ell(\theta_{t}+\sum_{j=1}^{k}\lambda_{j}(\tilde{s}_{j}-\theta_{t}),D)$.
@@
$g_{t}^{k}:=\max_{\lambda\in\Delta^{k}}\max_{s_{1},\ldots,s_{k}\in\mathcal{M}}\sum_{j=1}^{k}\lambda_{j}\langle\nabla\ell(\theta_{t},D),\theta_{t}-s_{j}\rangle.$
@@
$g_{t}=\max_{s\in\mathcal{M}}\langle\nabla\ell(\theta_{t},D),\theta_{t}-s\rangle,$
@@
because the top-$k$ FW gap subsumes the original FW gap by setting $\lambda_{1}=1$ and $\lambda_{j}=0$ for $j=2,\dots,k$. Intuitively, selecting multiple descent directions and optimizing their combination always gives at least as much descent as the single best direction. From the Lipschitz continuity of $\ell(\theta,D)$, we have:
@@
${\rm LMO}(\{\theta_{i}^{*,l}\}_{i=1}^{n},D,\theta_{t}^{l})=\operatorname*{arg\,min}_{s^{l}\in\{\theta_{1}^{*,l},\dots,\theta_{n}^{*,l}\}}\langle\nabla\ell(\theta^{t},D)^{l},s^{l}\rangle.$ (10)
@@ 6 Acknowledgement
We thank the Department of Applied Mathematics at the Hong Kong Polytechnic University for generously providing compute resources.

### v3

@@ Abstract
Model merging has emerged as a promising approach for multi-task learning (MTL), offering a data-efficient alternative to conventional fine-tuning. However, with the rapid development of the open-source AI ecosystem and the increasing availability of fine-tuned foundation models, existing model merging methods face two key limitations: (i) They are primarily designed for in-house fine-tuned models, making them less adaptable to diverse model sources with partially unknown model and task information, (ii) They struggle to scale effectively when merging numerous model checkpoints. To address these challenges, we formulate model merging as a constrained optimization problem and introduce a novel approach: Frank-Wolfe Merging (FW-Merging). Inspired by Frank-Wolfe optimization, our approach iteratively selects the most relevant model in the pool to minimize a linear approximation of the objective function and then executes a local merging similar to the Frank-Wolfe update. More importantly, FW-Merging serves as an orthogonal technique for existing merging methods, seamlessly integrating with them to further enhance accuracy performance. Our experiments show that FW-Merging scales across diverse model sources, remaining stable with 16 irrelevant models and improving by 15.3% with 16 relevant models on 20 CV tasks, while maintaining constant memory overhead—unlike the linear overhead of data-informed merging methods. Compared with the state-of-the-art approaches, FW-Merging surpasses the data-free merging method by 32.8% and outperforms the data-informed Adamerging by 8.39% when merging 20 ViT models. Our code is open-sourced at here.
@@ 1 Introduction
Figure 1: Performance scaling of FW-Merging across CV tasks. (a) demonstrate robustness to irrelevant models, while (b) show improved performance with relevant models. (c) analyzes performance degradation when incorporating a noisy model initialized from a different pre-trained checkpoint. Detailed results and experimental setup are discussed in Section $4.$
@@ 3.2 Frank-Wolfe Model Merging
We consider the problem of fine-tuning a pre-trained foundation model on new tasks. Given a pre-trained model $\theta_{0}$ and previously fine-tuned models $\{\theta_{1}^{*},\theta_{2}^{*},\cdots,\theta_{n}^{*}\}$, we aim to fine-tune the $(n+1)$-th model $\theta_{n+1}$ on new tasks as a convex combination of the previous models with optimal merging coefficients.
@@
$\min_{\lambda}\ell\Big(\sum_{i=1}^{n}\lambda_{i}\theta_{i}^{*}\Big)\quad\text{s.t.}\quad\sum_{i=1}^{n}\lambda_{i}=1,\quad\lambda_{i}\geq 0,$ (5)
where $\lambda=\{\lambda_{1},\ldots,\lambda_{n}\}$ are the merging coefficients, and $\ell$ is a loss function formulated for a specific goal, such as aligning the merged and individual models [63] or satisfying a task objective [69].
@@
$\min_{\theta\in\mathcal{M}}\ell(\theta),$ (6)
@@
$\min_{\theta\in\mathcal{M}}\ell(\theta)$ $=\min_{\lambda\in\Delta^{n}}\ell\Big(\sum_{i=1}^{n}\lambda_{i}\theta_{i}^{*}\Big).$
@@
${\rm LMO}\!\left(\{\theta_{i}^{*}\}_{i=1}^{n},\theta_{t}\right)=\operatorname*{arg\,min}_{s\in\rm\{\theta_{1}^{*},\ldots,\theta_{n}^{*}\}}\langle\nabla\ell(\theta_{t}),s\rangle$ (7)
@@
Input : Initial solution $\theta_{0}$; Fine-tuned checkpoints $\{\theta_{i}^{*}\}^{n}_{i=1}$; FW budget $T$.
@@
1:if $\theta_{0}\notin\mathcal{M}$ then $\mathcal{M}:=\mathcal{M}\cup\{\theta_{0}\}$
@@
4:  if $g_{t}:=\langle-\nabla\ell(\theta_{t}),d_{t}\rangle\leq\epsilon$ then
@@ 3.3 Design choices of the algorithm
In the case of deep learning, the optimization problem is non-convex, additional efforts are needed to better characterize the loss landscape and prevent the LMO from being dominated by one or a few fine-tuned models, which occurs because the linear approximation of $\ell(\theta)$ is an inaccurate sketch of the original objective function. Instead of relying on the argmin of linear subproblem, we fetch the top-$k$ vertices of LMO, $\{\tilde{s}_{j}\}_{j=1}^{k}$. A more subtle top-$k$ operation can be performed in a task-wise fashion if the original objective function involves multi-tasks.
@@
The merging function maintains feasibility, i.e., the merged model stays in the convex hull $\mathcal{M}$.
@@
Consider $\ell(\theta)$ be $L$-smooth over $\mathcal{M}$, which has two constants: ${\rm diam}:=\max_{\theta_{1},\theta_{2}\in\mathcal{M}}\|\theta_{1}-\theta_{2}\|$ be the diameter of $\mathcal{M}$, and ${\rm subopt}:=\ell(\theta_{0})-\min_{\theta\in\mathcal{M}}\ell(\theta)$ be the global suboptimality. Consider the soft FW algorithm which introduces the following changes to Algorithm 1:
@@
    $\{\lambda_{j}^{*}\}_{j=1}^{k}=\arg\min_{\lambda\in\Delta^{k}}\ell(\theta_{t}+\sum_{j=1}^{k}\lambda_{j}(\tilde{s}_{j}-\theta_{t}))$.
@@
$g_{t}^{k}:=\max_{\lambda\in\Delta^{k}}\max_{s_{1},\ldots,s_{k}\in\mathcal{M}}\sum_{j=1}^{k}\lambda_{j}\langle\nabla\ell(\theta_{t}),\theta_{t}-s_{j}\rangle.$
@@
$g_{t}=\max_{s\in\mathcal{M}}\langle\nabla\ell(\theta_{t}),\theta_{t}-s\rangle,$
@@
because the top-$k$ FW gap subsumes the original FW gap by setting $\lambda_{1}=1$ and $\lambda_{j}=0$ for $j=2,\dots,k$. Intuitively, selecting multiple descent directions and optimizing their combination always gives at least as much descent as the single best direction. From the Lipschitz continuity of $\ell(\theta)$, we have:
@@
${\rm LMO}(\{\theta_{i}^{*,l}\}_{i=1}^{n},D,\theta_{t}^{l})=\operatorname*{arg\,min}_{s^{l}\in\{\theta_{1}^{*,l},\dots,\theta_{n}^{*,l}\}}\langle\nabla\ell(\theta^{t})^{l},s^{l}\rangle.$ (10)
@@ 6 Acknowledgement
The support of the United Kingdom EPSRC (grant number UKRI256, EP/V028251/1, EP/N031768/1, EP/S030069/1, and EP/X036006/1), Intel, and AMD is gratefully acknowledged.